(x+5)/2+(2x-1)/5=5
I know the answer I just don't know how to get to it

Answer by algebrahouse.com(1659)

(x+5)/2 + (2x-1)/5 = 5
10[(x + 5)/2 + (2x - 1)/5 = 5] {multiplied entire equation by LCD, 10, to clear out fractions
10[(x + 5)/2] + 10[(2x - 1)/5] = 10(5) {multiplied each term by 10}
5(x + 5) + 2(2x - 1) = 50 {cancelled and multiplied what's left}
5x + 25 + 4x - 2 = 50 {used distributive property}
9x + 23 = 50 {combined like terms}
9x = 27 {subtracted 23 from both sides}
x = 3 {divided both sides by 9}
www.algebrahouse.com